OpenStudy (anonymous):

Hello there Im not sure what this problem is called and how to solve it so can you help me?

OpenStudy (anonymous):

\[\left| -3+1 \right|-\left| 5-2 \right|\]

OpenStudy (anonymous):

the answer is not -5

OpenStudy (anonymous):

2-3=-1